# Jianxiong Gao, Yichang Liu, Baofeng Yang, Jianfeng Feng, Yanwei Fu

### v1

@@ Appendix B EEG Experimental Device
We illustrate the electrode montage of a 64-channel EEG cap configured according to the GSN-HydroCel-64_1.0 layout in Fig. B.1.
@@
Figure B.1: Electrode montage of a 64-channel EEG cap using the GSN-HydroCel-64_1.0 layout. Sensor positions are annotated with their corresponding channel labels.
Figure B.2: More Results of CineSync: We present 12 frames with timestamps compared with the ground truth (GT).
Figure B.3: More Results of CineSync: We present 12 frames with timestamps compared with the ground truth (GT).
Figure B.4: More Results of CineSync: We present 12 frames with timestamps compared with the ground truth (GT).
@@
WER $\downarrow$ CER $\downarrow$ FAD $\downarrow$ LSD $\downarrow$ MCD $\downarrow$ SSIM $\uparrow$ RMSE $\downarrow$
CineSync-EEG 1.802 1.924 4.221 0.949 0.710 0.777 0.239
CineSync-fMRI 1.776 1.835 4.032 0.951 0.704 0.783 0.248
CineSync 1.476 1.559 3.811 0.927 0.683 0.814 0.221
CineSync 1.383 1.506 3.625 0.916 0.671 0.819 0.216
Table B.1: Performance comparison of CineSync with baselines. The average metrics across all subjects are reported. CineSync indicates the experiment that includes vision-related ROIs in fMRI. Bold denotes the best performance
Figure B.5: Qualitative results of CineSync and the two strong baselines are presented. “GT” indicates the ground-truth mel spectrogram.
Appendix C Prompts for Generating Text Descriptions
Textual descriptions are crucial for our training procedure. Prompts utilized for generating descriptions with Qwen2.5-VL [2] and Llava-Video [46] are presented in Fig. C.1 and Fig. C.2, respectively.
Figure C.1: Prompt utilized by Qwen2.5-VL for generating video descriptions.
Figure C.2: Prompt utilized by Llava-Video for generating video descriptions.
Appendix D Limitations and Future Work

### v2

@@ 7.3 EEG Experimental Device
We illustrate the electrode montage of a 64-channel EEG cap configured according to the GSN-HydroCel-64_1.0 layout in Fig. 8.
8 Additional Experimental Results
8.1 Implementation Details
[…]
Figure 7: Prompt utilized for generating video descriptions.
@@ 8.1 Implementation Details
Figure 8: Electrode montage of a 64-channel EEG cap using the GSN-HydroCel-64_1.0 layout. Sensor positions are annotated with their corresponding channel labels.
Table 4: Ablation study on multimodal alignment in CineSync. We report the average performance across all subjects. CineSync indicates the experiment that includes audio-related ROIs in fMRI.
@@
Video-based Frame-based Video-based Frame-based
[…]
w/o Vision 0.863 0.275 52.06 0.858 0.378 0.895 0.738 0.272 11.69
w/o Text 0.891 0.294 50.15 0.887 0.394 0.908 0.949 0.285 11.95
w/o Across 0.873 0.279 51.29 0.864 0.382 0.902 0.943 0.274 11.74
CineSync 0.926 0.336 44.77 0.954 0.423 0.921 0.953 0.297 12.18
Figure 9: Video Reconstruction Results for Subjects 1, 2, and 6. We compare the reconstructed frames from Subjects 1, 2, and 6 with the corresponding ground-truth (GT) frames. The consistent semantic alignment and visual fidelity across subjects demonstrate the robustness and strong cross-subject generalization ability of CineSync.
Figure 10: Video Reconstruction Results for Subjects 3, 4, and 5. Reconstructed frames from Subjects 3, 4, and 5 are shown alongside the GT frames at matched timestamps. The results further verify that CineSync maintains stable reconstruction quality across different individuals.
Table 5: Performance of each subject in CineBrain. CineSync indicates the experiment that includes audio-related ROIs in fMRI.
Methods Semantic-level Perceptual-level
Video-based Frame-based Video-based Frame-based
[…]
Subject 1 0.921 0.332 46.12 0.951 0.419 0.918 0.949 0.294 12.11
Subject 2 0.928 0.337 44.03 0.956 0.426 0.922 0.954 0.298 12.20
Subject 3 0.923 0.334 45.87 0.952 0.421 0.917 0.950 0.295 12.14
Subject 4 0.927 0.338 43.95 0.957 0.427 0.923 0.956 0.299 12.23
Subject 5 0.929 0.339 44.62 0.955 0.425 0.920 0.952 0.296 12.17
Subject 6 0.924 0.335 44.97 0.953 0.420 0.919 0.951 0.297 12.22
CineSync 0.926 0.336 44.77 0.954 0.423 0.921 0.953 0.297 12.18
Figure 11: More Results of CineSync: We present 12 frames with timestamps compared with the ground truth (GT).
Figure 12: More Results of CineSync: We present 12 frames with timestamps compared with the ground truth (GT).
Figure 13: More Results of CineSync: We present 12 frames with timestamps compared with the ground truth (GT).
8.2 Ablation Study on Multimodal Alignment
To quantify the contribution of each alignment component in CineSync, we conduct ablations on the full model trained with both audio- and vision-related ROIs in fMRI. Specifically, w/o Vision, w/o Text, and w/o Across correspond to removing (i) the vision–fMRI alignment branch, (ii) the text–fMRI alignment branch, and (iii) the cross-modal alignment between EEG and fMRI, respectively. The averaged results across all subjects are reported in Tab. 4. Removing any of these alignment pathways leads to clear and consistent performance drops across both semantic- and perceptual-level metrics, demonstrating that rich cross-modal alignment is critical for achieving high-quality brain-to-video reconstruction.
8.3 Detailed Results for Each Subject
To further verify that our model remains robust across individuals, we report per-subject quantitative performance in Tab. 5. As shown, the results exhibit only minor variations across the six subjects, indicating that CineSync generalizes well to different brain patterns. Moreover, the averaged scores closely match the overall CineSync results, further confirming subject-invariant performance.
8.4 Video Reconstruction Across Subjects
To better illustrate the robustness of our model, we visualize the reconstructed videos from different subjects under the same visual stimuli. Representative results are shown in Fig. 9 and Fig. 10. Since Subjects 1, 2, and 6 share the same train–test split, and Subjects 3, 4, and 5 share another split, we compare the reconstruction quality within each group accordingly. Specifically, Fig. 9 presents the reconstructions of Subjects 1, 2, and 6, while Fig. 10 shows those from Subjects 3, 4, and 5. These results indicate that our model achieves stable performance across different individuals.
8.5 More Results of Video Reconstruction
To further showcase the quality, semantic fidelity, and temporal coherence of the reconstructed videos, we present additional 12-frame examples in Fig. 11, Fig. 12, and Fig. 13. Across diverse scenes, our method consistently produces reconstructions that are semantically aligned with the stimuli and temporally smooth.
9 Limitations and Future Work